OpenStudy (drummmlers):

The hypotenuse of a right triangle is 10 inches. The shorter leg is 10 inches less than twice the longer leg. Find the lengths of the two legs.

OpenStudy (anonymous):

Do u have any choices to choose from?

OpenStudy (drummmlers):

No.

OpenStudy (anonymous):

okay then IDK..its easier when I have choices..sorry.

OpenStudy (anonymous):



OpenStudy (anonymous):

thats the longest side what twice of 10

OpenStudy (anonymous):

The long side is 10 and the short side is 10... i think

OpenStudy (drummmlers):

I have to solve it using a quadratic equation I'm pretty sure

OpenStudy (anonymous):

Hypotenuse= 10 Long side= x Short side = 2x-10 x= 2x -10 -x -x 0= x -10 +10 +10 10=x 10=2(10)-10

OpenStudy (drummmlers):

I have the answer but I have to show my work. The answers are 6 and 8

OpenStudy (anonymous):

oh ok nevermind... sorry